Show that the test-taking the overall significance of regression model using ANOVA table to be expressed as 𝑭=R2/𝒌−𝟏⁄(𝟏−𝑹𝟐)𝒏−𝒌
1
Expert's answer
2021-08-31T11:09:14-0400

"F=\frac{MSR}{MSE}"


Where;

"MSR= \frac{SSR}{1}"


And

"MSE = \frac{SSE}{n-2}"


"F=\frac{1-R^2}{n-k}"